OpenStudy (anonymous):

Which of the following linear equations are written in Standard Form? (Select all that apply.) Question 14 options: a) y=34x+1 b) 3x−5y=2 c) yx=59 d) 12x + 13y = 3 e) 3x+2y=18 f) 9y=5x ILL FAN

OpenStudy (anonymous):

@isaidavila

OpenStudy (mathmale):

First of all, what does "equation of a straight line in standard form" mean to you? We have to know what our goals look like before we can consistently achieve them;

OpenStudy (anonymous):

uhh idk

OpenStudy (isaidavila):

The standard equation is y=mx+b Just compare in the answers

563blackghost (563blackghost):

That is slope-intercept equation @isaidavila ... The standard form equation is c=ax+bx....

563blackghost (563blackghost):

meant c=ax+by...

OpenStudy (anonymous):

so c. or no and there can be more than one?

563blackghost (563blackghost):

Well it clearly states "Select all that apply." so yes there can be more than one ^^

OpenStudy (anonymous):

soory :P

563blackghost (563blackghost):

I dont think so for C...@mathmale c wont apply right?

OpenStudy (anonymous):

so was I right is it c

563blackghost (563blackghost):

@mathmale

OpenStudy (isaidavila):

Oh true thanks @563blackghost to remember me

OpenStudy (anonymous):

@isaidavila @563blackghost

OpenStudy (anonymous):

I don't understand

563blackghost (563blackghost):

dont know about c but what others would apply?

OpenStudy (anonymous):

b and e

OpenStudy (anonymous):

@563blackghost so b and e?

OpenStudy (anonymous):

@isaidavila is it b and e?

563blackghost (563blackghost):

Correct ^^ they do apply

563blackghost (563blackghost):

there is another one though...

OpenStudy (isaidavila):

Yes

OpenStudy (anonymous):

f?

OpenStudy (mathmale):

It's important that you can JUSTIFY

OpenStudy (isaidavila):

D also applies

OpenStudy (anonymous):

oh ok

563blackghost (563blackghost):

f does not apply...

OpenStudy (anonymous):

D applies

OpenStudy (mathmale):

... JUSTIFY your choice of answer. I will spill the beans and agree that C does NOT apply; that equation is NOT in standard form. Again I urge you to determine first what an equation in standard form of a straight line looks like.

OpenStudy (mathmale):

Here, for your reference, are several equations of straight lines: \[1) y=mx+b\]

OpenStudy (anonymous):

so b, d, and e ?

OpenStudy (mathmale):

\[2. y-y _{1}=m(x-x _{1})\]